# K-Act2Emo: Korean Commonsense Knowledge Graph for Indirect Emotional Expression

### 21 Mar 2024 v1

@@ 1 Introduction
Literary texts often convey characters’ emotions indirectly through descriptions of gestures, facial expressions, or appearances. For example, we can infer someone’s anxiety from their hand gestures. These nonverbal cues act as common knowledge within specific societies or cultures (Matsumoto and Hwang, 2016), allowing readers to interpret the emotional state of characters. However, datasets addressing this task have been confined to subsets of commonsense knowledge graphs (CSKGs) (Sap et al., 2019; Hwang et al., 2021). These often intertwine inferences on emotions with non-emotional elements, and their coverage is limited.
@@
이를 간다 Grinds one’s teeth NegEnv 분노하다 Angry
Table 1: Example Triple Sets from K-Act2Emo. English translations are based on the closest meanings, though the usage contexts for words marked with * do not fully align.
@@ 2 Background
ATOMIC (Sap et al., 2019) distinguishes itself from other Knowledge Graphs (KGs) such as CONCEPTNET (v5.7) (Speer et al., 2018) and TRANSOMCS (Zhang et al., 2020) through its specialized focus on event-based inferences and social commonsense knowledge. This encompasses dynamic event features like causes and effects, if-then relationships, and mental states. It is structured with 877K triples in the Head-Relation-Tail format, spanning 9 reasoning types (Relations), each meticulously collected and validated via crowdsourcing. The subsequent extension, $\mathrm{ATOMIC}^{20}_{20}$, broadens the scope with 1.33M triples across 23 reasoning types. K-Act2Emo adopts this triple format as well, but with a modification in the reasoning types to better accommodate inferences related to indirect emotional expressions.
@@
Table 2: Potentially related reasoning types in $\mathrm{ATOMIC}^{20}_{20}$. PersonX refers to the agent of the Head event
@@
KGs, while beneficial, grapple with issues of scalability and expressiveness, lacking the flexibility inherent in Language Models (LMs). To address this gap, there is an increasing interest in research that combines KGs with LMs and explores the use of LMs as dynamic knowledge bases. COMET ((Bosselut et al., 2019)) demonstrates that pre-trained LMs can leverage commonsense knowledge graphs to generate new, valid inferences about events. West et al. (2021) propose Symbolic Knowledge Distillation, a machine-to-corpus-to-machine approach that employs large language models to augment commonsense knowledge graphs. This method trains smaller-sized language models on these enhanced commonsense knowledge graphs to produce high-quality causal commonsense knowledge. In this study, we also aim to assess transformer-based LMs’ knowledge in emotional inference and the benefits of incorporating K-Act2Emo with LMs.
@@ 3.1 Collection
Indirect Emotional Expression Collection Our initial step in collecting indirect emotional cues involves crowdsourcing from 10 Korean native speakers who have a wealth of experience in creating literary works. We provide them with emotion-related words from Hong and Jeong (2009) and a comprehensive list of body parts (see Appendix A). To avoid repetitive contributions, we make the growing corpus publicly available in real-time, rewarding only those submissions that are genuinely original compared to the existing corpus. We consider actions as new if they exhibit significant differences in nuance. As generating novel expressions becomes increasingly difficult, we escalate the rewards and introduce guidelines to prevent direct emotional expressions, such as ’wearing a sad smile.’ After the collection phase, we standardize tenses and unify personal references (e.g., friends, family, lovers) under the singular term ’Y’ to ensure consistency.
@@
Reasoning Type Design From the insights gained through the two pilot tests, we design the reasoning types. The first phase, without distinguishing between reasoning types, identified a wide range of emotions across various scenarios, often skewed by the possibility of multiple interpretations. To mitigate this, we classify situations into positive (PosEnv) or negative (NegEnv) categories, offering clear examples like "receiving support" or "a family member passing away" to direct participants towards more precise contexts. The second pilot showed improved emotional inference generation; however, there were still issues with reporting non-emotional aspects, such as "being sick." To address these discrepancies, we introduced a new category, NonEmo Relation, to more accurately document these inferences.
Inference Collection In order to gather emotional expressions inferred from indirect cues, we conducted another crowdsourcing session with 20 Korean speakers. To grasp nuanced emotions and expedite the process, we allowed participants to freely describe emotions using adjectives and phrases, rather than selecting from fixed emotional categories. We encouraged them to use direct emotional words over metaphors, limiting descriptions to two words. Using the template "___ deut (as if he/she feels ___)" along with the situation descriptions for Relation (see Figure 2), participants filled in natural emotional expressions. Up to five inferences per indirect cue were permitted. If inferred emotions for both PosEnv and NegEnv contexts are identical, participants may select one. To avoid unnatural inferences, an option for No Inferred Emotion is provided. Participants are free to enter any inferences unrelated to emotional expressions in the NonEmo field.
@@ 3.2 Statistics
Similar to ATOMIC, the annotations in K-Act2Emo is a range of potential inferences. Therefore, to validate our dataset, we enlist 20 workers who were involved in the collection of inferences to indicate whether they agree with the contributions of other workers. In this process, we utilize a 4-point Likert scale (always/often, sometimes/likely, farfetched/never, invalid), used by Hwang et al. (2021), where the first two labels are considered Accept and the last two Reject. However, unlike Hwang et al. (2021), we did not provide a No judgement option. The evaluation results are as follows: out of a total of 7,272 nodes, 6,002 were accepted, resulting in an acceptance rate of 82.53%. The acceptance rate for nodes in NegEnv is slightly lower compared to that in PosEnv (Table 3).
@@
rel.NoneEmo 894 6 99.33 900
Table 3: Percentage of node acceptance in raw annotations evaluated by crowdsourcing participants.
@@
Node Analysis by Relation Table 4 details our knowledge graph, containing 6,002 nodes for 1,900 unique indirect emotions. It notes 634 nodes without positive context inferences and 135 without negative context inferences. Additionally, it details Tail distribution: 1,951 in PosEnv, 3,157 in NegEnv, and 894 in NonEmo, showing more inferences in NegEnv. There are 64, 94, and 124 unique Tails in PosEnv, NegEnv, and NonEmo, respectively, highlighting greater diversity in negative contexts. Except for "kancelhata" (Desperate), "kincanghata" (Nervous), "nollata" (Surprised), "tanghwanghata" (Embarrassed), and "pwukkulepta" (Ashamed), there is no overlap between the Tails used in PosEnv and NegEnv. The terms "mwukilyekhata" (Helpless), "casinkami issta" (Confident), and "hokisimi sayngkita" (Curious) are observed in both non-emotional and emotional Tails.
@@
Table 5: Analysis of word and morpheme counts in Head and Tail
@@
Table 6: Emotion category distribution across inference nodes in PosEnv and NegEnv Relations.
Number of Emotions Inferred per Head While the collection of responses in a free-form manner allows for the capture of a wide range of emotional expressions, categorizing these expressions is essential for identifying similar emotional states. Therefore, we employed the Korean emotional adjective and verb dictionary by Hong and Jeong (2009) to classify the inferred emotions into 21 distinct categories.
Table 6 presents the distribution of emotion categories across emotional inference nodes. In the PosEnv Relation, the categorization process resulted in a consolidation to 1,595 nodes across 12 Emotion Categories. In contrast, in the NegEnv Relation, this process led to a reduction to 2,659 nodes, encompassing 14 Emotion Categories. Emotion Categories common to both Relations include Surprise, Sympathy, Fear, Complex Emotion, Shame, and Desire.
Table 7 demonstrates the impact of categorization on the number of emotions inferred per Head. Initially, for Heads with three Tails, the counts were 148 in positive contexts and 361 in negative contexts. Post-categorization, these numbers reduced to 35 and 119 in positive and negative contexts, respectively. Similarly, for Heads with two Tails, counts decreased from 389 to 259 in positive contexts and from 670 to 656 in negative contexts. A significant observation is the predominance of Heads with three emotion categories in NegEnv compared to PosEnv. This phenomenon can be attributed to the higher arousal levels in negative contexts, which allow for simultaneous interpretations of Anger and Fear, contrasting with the distinct states of Joy and Relaxation in positive contexts.
@@
Table 7: Comparison of Head Count Based on Number of Tails/Emotion Categories in PosEnv and NegEnv
@@ 4 Knowledge Coverage Comparison
In this study, we compare K-Act2Emo with $\mathrm{ATOMIC}^{20}_{20}$. We translated K-Act2Emo’s indirect emotional expressions into English using DeepL [^3] https://www.deepl.com/translator, standardizing personal references to "PersonX" and replacing "Y" with "PersonY" for consistency.
@@
Table 9 displays Heads unique to K-Act2Emo as well as those found in both K-Act2Emo and $\mathrm{ATOMIC}^{20}{20}$. Despite $\mathrm{ATOMIC}^{20}{20}$’s extensive coverage of daily events, it often fails to include common expressions that convey emotions. This omission stems not merely from cultural variances but from its lack of specialization in indirect emotional expression. This underscores K-Act2Emo’s depth in understanding emotional expressions, distinguishing it from $\mathrm{ATOMIC}^{20}_{20}$.
@@
5 Methods
Experiment Setup In this study, we assess whether K-Act2Emo can be utilized for training emotional inference knowledge in Transformer-based models. We implement a hold-out evaluation strategy on the K-Act2Emo dataset, dividing nodes related to PosEnv and NegEnv into training, validation, and testing partitions comprising 4,124, 482, and 502 nodes, respectively. These subsets include 1,520, 180, and 180 unique heads, ensuring no overlap of heads between them. We also examine the distribution of emotion categories across these subsets to avoid bias (see Appendix  C for more details).
We train COMET-BART, a BART-based (Lewis et al., 2020) model using the COMET framework (Bosselut et al., 2019), with the K-Act2Emo training set to generate Tails given its Head and Relation. We use KoBART [^4] https://github.com/SKT-AI/KoBART as the base model. As a baseline, we compare it to a one-shot GPT-4 Turbo (0125-preview), using a single randomly sampled fact from the training dataset (with the same Relation as the testing fact). To further examine the emotional inference capabilities of prominent Korean LLMs, we conduct one-shot and five-shot learning with Mi:dm (KT, 2023), ChatSKKU [^5] https://huggingface.co/jojo0217/ChatSKKU5.8B, Kullm (Lee et al., 2023), KoAlpaca  [^6] https://huggingface.co/beomi/KoAlpaca-Polyglot-5.8B, and KoPlatypus [^7] https://huggingface.co/kyujinpy/KoT-platypus2-7B, using facts sampled from the training dataset for each. We choose one-shot over zero-shot learning because the performance in zero-shot learning is significantly lower, and it generates overly lengthy results compared to K-Act2Emo’s Tails. For details on each training session, refer to Appendix  B.
Evaluation Setup The evaluation of these models includes both automatic and human assessments. For automatic evaluations, we utilize BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), and BERT Score (Zhang et al., 2019) metrics. We used the Mecab-ko morpheme analyzer as a tokenizer for evaluation, and for the BERT Score, we utilized the KoBERT Score [^8] https://github.com/lovit/KoBERTScore.git based on the Kcbert (Lee, 2020) model. Human evaluations were conducted by five reviewers on the outputs from COMET-BART and five-shot Mi:dm. They evaluated the acceptability of these models based on the criteria established by Hwang et al. (2021).
6 Results
We present our principal findings in Table 10, Table 11, and Table 12. Table 10 demonstrates that COMET-BART, trained with K-Act2Emo, outperforms other models across all automatic metrics. A significant improvement is observed in all models when employing 5-shot learning over 1-shot learning. Besides COMET-BART, GPT-4 Turbo and Mi:dm emerge as the top performers. As detailed in Table 11, these models demonstrate a high degree of conformity to the response format of the provided demonstrations. Remarkably, despite Mi:dm’s low performance in zero-shot learning trials in the pilot test, it shows rapid adaptation after a single example. Table 11 illustrates an improvement in the response format with 5-shot learning compared to 1-shot learning across most of the models.
@@ 6 Results
In human evaluations, consistent performance differences are observed (see Table 12). GPT-4 turbo one-shot and Mi:dm 5-shot performances are rated 1.8 percentage points and 5.18 percentage points lower, respectively, compared to COMET-BART. These disparities are attributed to the inclusion of content related to characters’ personalities and states rather than emotions (see Table 13). Additionally, some grammatical errors are observed in the outcomes of Mi:dm. Although NonEmo-related nodes are not utilized in the experiments, these results suggest a necessity for leveraging such nodes to enhance alignment and reranking of generated responses.
@@
Tenses muscles Strong
@@
Participates in the conversation Listens attentively
@@
Tears at clothes Starving
@@
Mouth is slightly open Happy and expectant*
@@
Rubs lips together Feels cold
@@
Raises hand Asks for help

### 23 Mar 2024 v2

@@ 1 Introduction
Literary texts often convey characters’ emotions indirectly through descriptions of gestures, facial expressions, or appearances. For example, we can infer someone’s anxiety from their hand gestures. These nonverbal cues act as common knowledge within specific societies or cultures (Matsumoto and Hwang, 2016), allowing readers to interpret the emotional state of characters. However, datasets addressing this task have been confined to subsets of commonsense knowledge graphs (CSKGs) (Speer et al., 2018;Sap et al., 2019; Zhang et al., 2020, Hwang et al., 2021). These often intertwine inferences on emotions with non-emotional elements, and their coverage is limited.
@@
이를 간다 Grind one’s teeth NegEnv 분노하다 Angry
Table 1: Example triple sets from K-Act2Emo. English translations are based on the closest meanings, though the usage contexts for words marked with an asterisk do not fully align.
@@ 2 Background
ATOMIC (Sap et al., 2019) distinguishes itself from other Knowledge Graphs (KGs) such as CONCEPTNET (v5.7) (Speer et al., 2018) and TRANSOMCS (Zhang et al., 2020) through its specialized focus on event-based inferences and social commonsense knowledge. This encompasses dynamic event features like causes and effects, if-then relationships, and mental states. It is structured with 877K triples in the Head-Relation-Tail format, across 9 reasoning types (Relations), each collected and validated via crowdsourcing. The subsequent extension, $\mathrm{ATOMIC}^{20}_{20}$, broadens the scope with 1.33M triples across 23 reasoning types. K-Act2Emo adopts this triple format as well, but with a modification in the reasoning types to better accommodate inferences from indirect emotional expressions.
@@
Table 2: Reasoning types in $\mathrm{ATOMIC}^{20}_{20}$ which are potentially related to emotions. PersonX refers to the agent of the Head event.
@@
KGs, while beneficial, grapple with issues of scalability and expressiveness, lacking the flexibility inherent in Language Models (LMs). To address this gap, there is an increasing interest in research that combines KGs with LMs and explores the use of LMs as dynamic knowledge bases. COMET (Bosselut et al., 2019) demonstrates that pre-trained LMs can leverage commonsense knowledge graphs to generate new, valid inferences about events. West et al. (2021) propose Symbolic Knowledge Distillation, a machine-to-corpus-to-machine approach that employs large language models to augment commonsense knowledge graphs. This method trains smaller-sized language models on these enhanced commonsense knowledge graphs to produce high-quality causal commonsense knowledge. In this study, we also aim to assess transformer-based LMs’ knowledge in emotional inference and the benefits of incorporating K-Act2Emo with LMs.
@@ 3.1 Collection
Indirect Emotional Expression Collection Our initial step in collecting indirect emotional cues involved crowdsourcing from 10 Korean native speakers with extensive experience in creating literary works. We provided them with emotion-related words from Hong and Jeong (2009) and a comprehensive list of body parts (see Appendix A). To avoid repetitive submissions, we made the growing corpus publicly available in real-time, rewarding only those submissions that were genuinely original compared to the existing corpus. We considered indirect emotional expressions as new if they exhibited significant differences in nuance. As generating novel expressions became increasingly difficult, we escalated the rewards and introduced guidelines to prevent direct emotional expressions, such as "wearing a sad smile." After the collection phase, we standardized verb tenses to present tense and unified personal references (e.g., friends, family, lovers) under the singular term "Y" to ensure consistency.
@@
Reasoning Type Design Before gathering emotions from indirect expressions, we conducted two pilot tests to design proper reasoning types. The first phase, when we collected emotions without reasoning types, identified a wide range of emotions across various scenarios, often skewed by the possibility of multiple interpretations. To mitigate this, we classified situations into positive (PosEnv) or negative (NegEnv) categories, offering clear examples like "receiving support" or "a family member passing away" to direct participants towards more precise contexts. The second pilot, which involved collecting emotions with both positive and negative situational factors (PosEnv and NegEnv), resulted in consistent emotional inferences within an event; however, there were still issues with reporting non-emotional aspects, such as "being sick." To address these discrepancies, we introduced a new category, NonEmo Relation, to more accurately document these inferences.
Inference Collection In order to gather emotional expressions inferred from indirect cues, we conducted another crowdsourcing session with 20 Korean speakers. To grasp nuanced emotions and expedite the process, we allowed participants to freely describe emotions using adjectives and phrases, rather than selecting from fixed emotional categories. We encouraged them to use direct emotional words over metaphors, limiting descriptions at most three words. Using the template "___ deut (as if he/she feels ___)" along with the situation descriptions for Relation (see Figure 2), participants filled in natural emotional expressions. Up to five inferences per indirect cue were permitted. If emotions inferred for both PosEnv and NegEnv contexts were identical, participants could select one. To avoid unnatural inferences, an option for No Inferred Emotion was provided. Participants were free to enter any inferences unrelated to emotional expressions in the NonEmo field.
@@ 3.2 Statistics
Similar to ATOMIC, the annotations in K-Act2Emo is a range of potential inferences. Therefore, to validate our dataset, we enlist 20 workers who were involved in the collection of inferences to indicate whether they agree with the contributions of other workers. In this process, we utilize a 4-point Likert scale (always/often, sometimes/likely, farfetched/never, invalid), used by Hwang et al. (2021), where the first two labels are considered Accept and the last two Reject. However, unlike Hwang et al. (2021), we do not provide a No judgement option. The evaluation results are as follows: out of a total of 7,272 nodes, 6,002 were accepted, resulting in an acceptance rate of 82.53%. The acceptance rate for nodes in NegEnv is slightly lower compared to that in PosEnv (Table 3).
@@
rel.NonEmo 894 6 99.33 900
Table 3: Percentage of node acceptance in raw annotations evaluated by crowdsourcing participants
Node Analysis by Relation Table 4 details our knowledge graph, containing 6,002 nodes for 1,900 unique indirect emotions. It notes 634 nodes without positive context inferences and 135 without negative context inferences. Additionally, it details Tail distribution: 1,951 in PosEnv, 3,157 in NegEnv, and 894 in NonEmo, showing more inferences in NegEnv. There are 64, 94, and 124 unique Tails in PosEnv, NegEnv, and NonEmo, respectively, highlighting greater diversity in negative contexts. Except for "kancelhata" (Desperate), "kincanghata" (Nervous), "nollata" (Surprised), "tanghwanghata" (Embarrassed), and "pwukkulepta" (Shy/Ashamed), there is no overlap between the Tails used in PosEnv and NegEnv. The terms "mwukilyekhata" (Helpless), "casinkami issta" (Confident), and "hokisimi sayngkita" (Curious) are observed in both non-emotional and emotional Tails.
@@
Table 5: Analysis of word and morpheme counts in Head and Tail.
@@
Table 6: Emotion category distribution across inference nodes in PosEnv and NegEnv Relations
Number of Emotions Inferred per Head While the collection of responses in a free-form manner allows for the capture of a wide range of emotional expressions, categorizing these expressions is essential for identifying similar emotional states. Therefore, we employ the Korean emotional adjective and verb dictionary by Hong and Jeong (2009) to classify the inferred emotions into 21 distinct categories.
Table 6 presents the distribution of emotion categories across emotional inference nodes. In the PosEnv Relation, the categorization process resulted in a consolidation to 1,595 nodes across 12 Emotion Categories. In contrast, in the NegEnv Relation, this process led to a reduction to 2,659 nodes, encompassing 14 Emotion Categories. Emotion Categories common to both Relations include Surprise, Sympathy, Fear, Complex Emotion, Shame, and Desire. [^3] Fear and Shame appear in PosEnv nodes because ”pwukkulepta” (Shy/Ashamed) is classified as Shame, and ”kincanghata” (Nervous) (with a neutral nuance in Korean) is classified as Fear.
The number of Tails per Head is significant as it can indicate the complexity of interpreting the Head. Table 7 demonstrates this complexity through Head counts per the number of Tails and emotion categories. For Heads with three Tails, the counts were 148 in PosEnv and 361 in NegEnv. After categorizing Tails into emotion categories, these numbers reduced to 35 and 119 in PosEnv and NegEnv, respectively. Similarly, for Heads with two Tails, counts decreased from 389 to 259 in PosEnv and from 670 to 656 in NegEnv. This reduction may be due to oversimplification because of categorization, but it also suggests that several Heads have Tails with similar meanings. A significant observation is the predominance of Heads with three emotion categories in NegEnv compared to PosEnv. This phenomenon can be attributed to the higher arousal levels in negative contexts, which allow for simultaneous interpretations of Anger and Fear, contrasting with the distinct states of Joy and Relaxation in positive contexts.
@@
Table 7: Comparison of Head count based on number of Tails/Emotion categories in PosEnv and NegEnv.
@@ 4 Knowledge Coverage Comparison
In this study, we compare K-Act2Emo with $\mathrm{ATOMIC}^{20}_{20}$. We translate K-Act2Emo’s indirect emotional expressions into English using DeepL [^4] https://www.deepl.com/translator, standardizing personal references to "PersonX" and replacing "Y" with "PersonY" for consistency.
@@
Table 9 displays Heads unique to K-Act2Emo as well as those found in both K-Act2Emo and $\mathrm{ATOMIC}^{20}_{20}$. Despite $\mathrm{ATOMIC}^{20}_{20}$’s extensive coverage of daily events, it often fails to include common expressions that convey emotions. This omission stems not merely from cultural variances but from its lack of specialization in indirect emotional expression. This underscores K-Act2Emo’s depth in understanding emotional expressions, distinguishing it from $\mathrm{ATOMIC}^{20}_{20}$.
@@
KG Type Instance
K-Act2Emo Head PersonX looks around slowly.
PosEnv peaceful, relaxed, tense
NegEnv aware, tense, nervous
NonEmo cautious
$\mathrm{ATOMIC}^{20}_{20}$ Head PersonX glances around the room.
xReact aware, attentive, tired, relieved, relaxed
xAttribute aware, concerned, cautious, curious, etc.
Even when addressing similar Heads and Tails, K-Act2Emo’s reasoning taxonomy proves to be more effective than that of $\mathrm{ATOMIC}^{20}_{20}$. As demonstrated in Table 10, K-Act2Emo distinctly categorizes content into PosEnv, NegEnv, and NonEmo. In contrast, $\mathrm{ATOMIC}^{20}_{20}$ not only redundantly repeats information across xAttr and xReact but also merges contents of disparate characteristics, complicating the analysis and application of inferences. This distinction highlights K-Act2Emo’s refined taxonomy in managing emotional expressions, offering a more structured framework for interpreting complex emotional cues.
@@
5 Emotional Inference Experiment
Experiment Setup In this study, we assess whether K-Act2Emo can be utilized for training emotional inference knowledge in Transformer-based models. We implement a hold-out evaluation strategy on the K-Act2Emo dataset, dividing nodes related to PosEnv and NegEnv into training, validation, and testing partitions comprising 4,124, 482, and 502 nodes respectively. These subsets include 1,520, 180, and 180 unique heads, ensuring no overlap of heads between them. We also examine the distribution of emotion categories across these subsets to avoid bias (see Appendix  C for more details).
We fine-tune COMET-BART, a BART-based (Lewis et al., 2020) model using the COMET framework (Bosselut et al., 2019), with the K-Act2Emo training set to generate Tails given its Head and Relation. We use KoBART [^5] https://github.com/SKT-AI/KoBART as the base model. As a baseline, we compare it to a one-shot GPT-4 Turbo (0125-preview), using a single randomly sampled fact from the training dataset (with the same Relation as the testing fact). To further examine the emotional inference capabilities of prominent Korean LLMs, we conduct one-shot and five-shot learning with Mi:dm (KT, 2023), ChatSKKU [^6] https://huggingface.co/jojo0217/ChatSKKU5.8B, Kullm (Lee et al., 2023), KoAlpaca  [^7] https://huggingface.co/beomi/KoAlpaca-Polyglot-5.8B, and KoPlatypus [^8] https://huggingface.co/kyujinpy/KoT-platypus2-7B, using facts sampled from the training dataset for each. We choose one-shot over zero-shot learning because the performance in zero-shot learning is significantly lower, and it generates overly lengthy results compared to K-Act2Emo’s Tails. For details on each training session, refer to Appendix  B.
Evaluation Setup The evaluation of these models includes both automatic and human assessments. For automatic evaluations, we utilize BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), and BERT Score (Zhang et al., 2019) metrics, with the tails for the test set as gold labels. We use the Mecab-ko morpheme analyzer as a tokenizer for evaluation, and for the BERT Score, we utilize KoBERTScore [^9] https://github.com/lovit/KoBERTScore.git based on the Kcbert (Lee, 2020) model. Human evaluations were conducted by five reviewers on the outputs from COMET-BART, one-shot GPT-4 turbo, and five-shot Mi:dm. They evaluated the acceptability of these models based on the criteria established by Hwang et al. (2021).
@@ 5 Emotional Inference Experiment
Results We present our principal findings in Table 11, Table 12, and Table 13. Table 11 demonstrates that COMET-BART, trained with K-Act2Emo, outperforms other models across all automatic metrics. [^10] Due to the overall brevity of the target Tails, the results for BLEU-4 were low. A significant improvement is observed in all models when employing five-shot learning over one-shot learning. Besides COMET-BART, GPT-4 Turbo and Mi:dm emerge as the top performers. As detailed in Table 12, these models demonstrate a high degree of conformity to the response format of the provided demonstrations. Remarkably, despite Mi:dm’s low performance in zero-shot learning trials in the pilot test, it shows rapid adaptation after a single example. Table 12 illustrates an improvement in the response format with five-shot learning compared to one-shot learning across most of the models.
@@
In human evaluations, consistent performance differences are observed (see Table 13). GPT-4 turbo one-shot and Mi:dm five-shot performances are rated 1.8 percentage points and 5.18 percentage points lower, respectively, compared to COMET-BART. These disparities are attributed to the inclusion of content related to characters’ personalities and states rather than emotions (see Table 14). Additionally, several grammatical errors are observed in the outcomes of Mi:dm. Although NonEmo-related nodes are not utilized in the experiments, these results suggest a necessity for leveraging such nodes to enhance alignment and reranking of generated responses.
@@
Tense muscles Strong
@@
Participate in the conversation Listen attentively
@@
Tear clothes Starving
@@
One’s mouth is slightly open Happy and expectant*
@@
Rub his/her lips Feel cold
@@
Raise his/her hand Ask for help